# Gauge freedoms in unravelled quantum dynamics: When do different continuous measurements yield identical quantum trajectories?

### v1

@@ Abstract
Quantum trajectories of a Markovian open quantum system arise from the back-action of measurements performed in the environment with which the system interacts. In this work, we consider counting measurements of quantum jumps, corresponding to different representations of the same quantum master equation. We derive necessary and sufficient conditions under which these different measurements give rise to the same unravelled quantum master equation, which governs the dynamics of the probability distribution over pure conditional states of the system. Since that equation uniquely determines the stochastic dynamics of a conditional state, we also obtain necessary and sufficient conditions under which different measurements result in identical quantum trajectories. We then consider the joint stochastic dynamics for the conditional state and the measurement record. We formulate this in terms of labelled quantum trajectories, and derive necessary and sufficient conditions under which different representations lead to equivalent labelled quantum trajectories, up to permutations of labels. As those conditions are generally stricter, we finish by constructing coarse-grained measurement records, such that equivalence of the corresponding partially-labelled trajectories is guaranteed by equivalence of the trajectories alone. These general results are illustrated by two examples that demonstrate permutation of labels, and equivalence of different quantum trajectories.
I Introduction
@@
Open quantum systems are important in many physical contexts where the influence of external environments cannot be neglected [1, 2, 3, 4]. They are often analysed using the quantum master equation (QME) [5, 6] where the system state is described via a density matrix. This averaged state follows a non-unitary evolution, due to the interaction with the environment. In contrast to open classical systems, the QME does not uniquely prescribe stochastic dynamics of fluctuating trajectories of the system. In fact, a single QME permits many different unravellings, each corresponding to a specific stochastic process for a conditional system’s state. Such processes arises from the back-action of a chosen continuous measurement on the environment. These possibilities appear in the QME via a set of gauge freedoms: the different stochastic processes are associated to different representations of the quantum master operator, corresponding to different decompositions in terms of jump operators and a system Hamiltonian.
@@
The dynamics of quantum trajectories has gained increased interest in recent years, covering subjects such as measurement induced (and other) phase transitions [7, 8, 9, 10, 10, 11, 12, 13, 14, 15, 16, 17], quantum control [18, 19, 20, 21], quantum stochastic thermodynamics [22, 23] and steady state ensemble preparation [24, 25, 26]. Experimentally, quantum trajectories are obtained from continuous monitoring of the system, e.g., by detecting its output into the environment. Advances in experimental techniques have now produced practical platforms in which to investigate phenomena related to these stochastic trajectories [27, 28, 29, 30, 31, 32, 33, 34]. In each case, the representation, which is determined by the continuous monitoring procedure used, plays a crucial role and different choices can have drastic impacts on the dynamics, including phenomena such as transport and entanglement entropy, as well as classical simulability [35, 36, 37, 38, 39, 40]. Indeed, trajectories play an important role in the numerical simulation of quantum master equations [41, 42, 43, 35, 44, 45, 46], which exploit the reduced dimensionality of pure conditional states, as opposed to (generically) mixed density matrices.
@@
In this work, we associate each representation of the QME with a piecewise deterministic stochastic process (PDP) [1]. This encodes a probability distribution over quantum trajectories, which we call an ensemble of quantum trajectories. However, the relationship between representations and PDPs is not one-to-one: there are many representations that encode the same ensemble. As a simple example, multiplying any jump operator by a phase factor does not change either the QME or the PDP. Since the same process can be represented in many different ways, we refer to this as gauge invariance and to the corresponding transformation as a gauge freedom.
@@
This work characterises these gauge freedoms by establishing necessary and sufficient conditions under which two representations of the quantum master operator lead to the same PDP. This is achieved by analysing the generators of the PDPs, which constitute unravelled QMEs. In contrast to the QME, they govern the dynamics of probability distributions for pure density matrices, which represent conditional system states. In fact, they fully determine the corresponding PDPs, and thus the ensembles of quantum trajectories. Interestingly, the resulting gauge freedoms are much richer than simple multiplication of jump operators by phase factors or even their permutations; this is particularly apparent in systems with reset jumps, which correspond to jump operators of rank 1, meaning that each jump operator resets the system’s conditional state to a fixed destination.
@@
In the physical context, gauge invariance or equivalence of PDPs and quantum trajectory ensembles are relevant for weak symmetries of quantum master equations, and whether these are inherited by quantum trajectory ensembles. These issues are explored in [47], which relies extensively on the results presented here. When the goal is to generate quantum trajectory ensembles either in experiments or simulations, our work clarifies the remaining gauge freedom that can be used to optimise these setups, e.g., by implementing instead a minimal equivalent representation. Our results also reinforce the special status of reset jumps, which are common in many physical scenarios; for example energy level transitions in quantum optics experiments [48, 49, 50], superconducting circuits [51, 34] and quantum dots [52, 53]. They also appear in the context of many-body quantum systems, e.g., when global projective measurements are performed at fixed rate [54]. Finally, there has been much recent interest in a notion of ‘resetting’ of quantum systems [55, 56, 57], which when to be implemented at a constant rate can be encoded into the standard master equation formalism using necessarily global reset jumps.
The paper is structured as follows. In Sec. II we review the QME and unravelled QMEs which prescribe stochastic quantum trajectories corresponding to different representations of the former. In Sec. III we give the conditions for the unravelled QMEs for different representations to be the same and the resulting implications. Sec. IV discusses labelled and partially-labelled quantum trajectories, and the equivalence conditions for associated stochastic processes. Sec. V presents an illustrative example. Sections VI and VII detail the derivation of the results given in the sections III and IV respectively. We conclude in Sec. VIII.
II System State Dynamics
@@
II.1 Quantum master equation
@@
is called a representation of the quantum master operator (which includes $d$ jump operators in this case).
@@ II.2 Quantum trajectories
is the transition rate facilitated by jump operator of type $k$, i.e., the rate for jump of type $k$, from current state $\psi$ into state $\psi^{\prime}$. In Eq. (13), we introduced a gradient, denoted $\nabla$. The action of this gradient on scalar functions $f(\psi)$ gives the matrix with elements $(\nabla f)_{jk}=\partial f/\partial\psi_{jk}$. The divergence of a matrix $M(\psi)$ is therefore ${\nabla\cdot M=\sum_{jk}(\partial M_{jk}/\partial\psi_{jk})}$ [64].
@@
where $\rho_{t}$ is the average state of the system (in general a mixed state). Taking the time derivative of this equation and using Eq. (II.2) with $f(\psi)=\psi$ shows that $\rho_{t}$ indeed obeys the QME in Eq. (1). [^2] This $f(\psi)$ is a matrix valued function; in such cases $\mathcal{W}$ acts separately on each matrix element.
@@
III Gauge Invariance
for Quantum Trajectories
This Section formulates the gauge freedoms of the unravelled quantum dynamics. These gauge freedoms are found by identifying which representations of the QME lead to the same unravelled generator. The results are stated here, with proofs given Sec. VI.
@@ III.1 SJED
$k,k\in S_{\alpha}\quad\Leftrightarrow\quad\forall{|\psi\rangle}\;\;\exists\,{c,c^{\prime}}\;\;c\,J_{k}|\psi\rangle=c^{\prime}J_{k^{\prime}}|\psi\rangle,\;\;$ (17)
@@
From a physical perspective, reset SJEDs are interesting because rank-1 jump operators appear in many physical settings, as discussed in Sec. I. Furthermore, for dynamics with jump operators of this type only, a semi-Markov mapping of the unravelled dynamics exists [63, 64, 65], which simplifies the sampled space in stochastic simulations. Non-reset SJEDs can be relevant in experimental settings if the environment consists of multiple identical reservoirs or classical noise is present in the measurement process. In simulations it would be natural to exploit the associated gauge freedom and combine them into a single jump operator from the start.
III.2 Equality of unravelled generators
@@
the corresponding SJED are denoted as $S_{1},\dots,S_{d_{C}}$ and $\tilde{S}_{1},\dots,\tilde{S}_{\tilde{d}_{C}}$, and give rise to the associated super-operators for the composite action of their jump operators [cf. Eq. (18)]
@@
The numbers of SJED are $d_{C}\leq d$ and $\tilde{d}_{C}\leq\tilde{d}$. Then the conditions for these two representations to describe the same PDP are given by the following Theorem.
@@
$\tilde{d}_{C}=d_{C}\quad\text{and}\quad\tilde{\mathcal{A}}_{\alpha}=\mathcal{A}_{\pi(\alpha)}\;\;\forall\alpha,\;$ (28b)
for some permutation $\pi$ of $\{1,2,\dots,d_{C}\}$.
Note that the definition of SJED ensures that the permutation $\pi$ appearing in Theorem $1$ is uniquely defined (it is not possible that ${\cal A}_{\alpha}={\cal A}_{\beta}$ for $\alpha\neq\beta$). The fact that Eq. (28) is sufficient for Eq. (27) can be verified directly from the definition of ${\cal W}$, with the aid of Eq. (20), as we now explain. Showing the converse requires more work, this proof is given in Sec. VI.
@@
This case is considered explicitly in the proof in Sec. VI. In general, while Eq. (28b) constrains the composite action of SJEDs, it leaves more freedom in the choice of jump operators, compared with Eq. (29) (see Sec. III.3, below).
An important corollary of Theorem 1 is that $\mathcal{W}^{\dagger}=\tilde{\mathcal{W}}^{\dagger}$ under the conditions in Eq. (28), which means that their stationary distributions over conditional states are identical, and also $P(\psi,t)$ at any time $t$ is the same for both representations provided that one considers the same initial distribution. The two different representations actually produce the same ensemble of quantum trajectories in that case [but Eq. (28) does not guarantee equivalence of their measurement records, see Sec. IV]. We clarify what transformations between representations allow for that invariance next.
@@
III.3 Gauge freedoms of unravelled quantum dynamics
Given Theorem 1, a natural question follows: which representations give rise to a given a set of super-operators $\{{\cal A_{\alpha}}\}_{\alpha=1}^{d_{C}}$, and hence to the same unravelled dynamics? To answer this, we note that equality of $\tilde{\mathcal{A}}_{\alpha}$ and $\mathcal{A}_{\pi(\alpha)}$ in Eq. (28b) is the condition [cf. Eq. (18)]
$\sum_{j\in\tilde{S}_{\alpha}}\tilde{\mathcal{J}}_{j}=\sum_{k\in S_{\pi(\alpha)}}{\mathcal{J}}_{k}\,.$ (30)
From Eq. (8), it then follows that the sets $\{\tilde{J}_{j}\}_{j\in\tilde{S}_{\alpha}}$ and $\{J_{k}\}_{k\in S_{\pi(\alpha)}}$ are different representations of the same completely positive super-operator, $\tilde{\mathcal{A}}_{\alpha}=\mathcal{A}_{\pi(\alpha)}$, and thus are related by an isometry [59]. Appendix A.2 describes the gauge freedoms associated with the composite action operator for each SJED.
Using these results, we now describe the full gauge freedom of the unravelled quantum dynamics. [See Appendix A.3 for details]. It is convenient to consider a minimal representation (in the sense that each ${\cal A}_{\alpha}$ is represented by a minimal number of jump operators, see Appendix A). Then the gauge freedoms of the unravelled dynamics are given as follows:
Suppose that $H^{\prime},J_{1}^{\prime},\dots,J_{d^{\prime}}^{\prime}$ is a representation of a given quantum master operator in which all SJED actions ${\cal A}^{\prime}_{1},\dots,{\cal A}^{\prime}_{d_{C}^{\prime}}$ have minimal representations. Then $H,J_{1},\dots,J_{d}$ has the same unravelled generator if and only if $d_{C}^{\prime}=d_{C}$ and there exists a permutation $\pi$ of $\{1,\dots,d_{C}\}$ and a matrix $\mathbf{V}\in\mathbb{C}^{d\times d^{\prime}}$ such that
@@ III.3 Gauge freedoms of unravelled quantum dynamics
$\mathbf{V}_{jk}^{(\alpha)}=0\quad\text{unless}\quad j\in S_{\alpha},\,k\in S^{\prime}_{\pi(\alpha)}$ (31c)
@@
$\sum_{j\in S_{\alpha}}[\mathbf{V}^{(\alpha)}_{jk}]^{*}\mathbf{V}^{(\alpha)}_{jk^{\prime}}=\delta_{kk^{\prime}}\quad\text{for}\quad k,k^{\prime}\in S^{\prime}_{\pi(\alpha)}\;.$ (31d)

### v3

@@ Abstract
Quantum trajectories of a Markovian open quantum system arise from the back-action of measurements performed in the environment with which the system interacts. In this work, we consider counting measurements of quantum jumps, and the associated representations of the quantum master equation. We derive necessary and sufficient conditions under which different measurements give rise to the same unravelled quantum master equation, which governs the dynamics of the probability distribution over pure conditional states of the system. Since that equation uniquely determines the stochastic dynamics of a conditional state, we also obtain necessary and sufficient conditions under which different measurements result in identical quantum trajectories. We then consider the joint stochastic dynamics for the conditional state and the measurement record. We formulate this in terms of labelled quantum trajectories, and derive necessary and sufficient conditions under which different representations lead to equivalent labelled quantum trajectories, up to permutations of labels. As those conditions are generally stricter, we finish by constructing coarse-grained measurement records, such that equivalence of the corresponding partially-labelled trajectories is guaranteed by equivalence of the trajectories alone. These general results are illustrated by two examples that demonstrate permutation of labels, and equivalence of different quantum trajectories.
1 Introduction
@@
Motivation – Open quantum systems are important in many physical contexts where the influence of external environments cannot be neglected [1, 2, 3, 4]. They are often analysed using the quantum master equation (QME) [5, 6] where the system state is described via a density matrix. This averaged state follows a non-unitary evolution, due to the interaction with the environment. In contrast to open classical systems, the QME does not uniquely prescribe stochastic dynamics of fluctuating trajectories of the system. In fact, a single QME permits many different unravellings, each corresponding to a specific stochastic process for a conditional system’s state. Such processes arises from the back-action of a chosen continuous measurement on the environment. These possibilities appear in the QME via a set of gauge freedoms: the different stochastic processes are associated to different representations of the quantum master operator, corresponding to different decompositions in terms of jump operators and a system Hamiltonian.
@@
The dynamics of quantum trajectories has gained increased interest in recent years, covering subjects such as measurement induced (and other) phase transitions [7, 8, 9, 10, 10, 11, 12, 13, 14, 15, 16, 17], quantum control [18, 19, 20, 21], quantum stochastic thermodynamics [22, 23, 24] and steady state ensemble preparation [25, 26, 27]. Experimentally, quantum trajectories are obtained from continuous monitoring of the system, e.g., by detecting its output into the environment. Advances in experimental techniques have now produced practical platforms in which to investigate phenomena related to these stochastic trajectories [28, 29, 30, 31, 32, 33, 34, 35]. In each case, the representation, which is determined by the continuous monitoring procedure used, plays a crucial role and different choices can have drastic impacts on the dynamics, including phenomena such as transport and entanglement entropy, as well as classical simulability [36, 37, 38, 39, 40, 41]. Indeed, trajectories play an important role in the numerical simulation of quantum master equations [42, 43, 44, 36, 45, 46, 47, 48], which exploit the reduced dimensionality of pure conditional states, as opposed to (generically) mixed density matrices.
With the focus of applications on quantum trajectories and their properties, important questions remain about correspondence between their dynamics and continuous measurements or unravellings. In particular, while the choice of continuous measurement uniquely defines the quantum trajectory dynamics, the converse does not hold. This non-uniqueness is the very freedom in the choice of experimental and numerical protocols for the generation of quantum trajectories.
@@
Theoretical framework – Here, we consider each representation of the QME, and its stochastic Schrödinger equation [49, 50, 2], which generates a piecewise-deterministic stochastic process (PDP) [1]. This encodes a probability distribution over quantum trajectories, which we call an ensemble of quantum trajectories. However, the relationship between representations and PDPs is not one-to-one: there are many representations that encode the same ensemble. As a simple example, multiplying any jump operator by a phase factor does not change either the QME or the PDP. Since the same process can be represented in many different ways, we refer to this as gauge invariance of quantum trajectories and to the corresponding transformation as a gauge freedom. In contrast to the well-established gauge freedoms of the QME [51], gauge invariance of quantum trajectory dynamics has not been characterised before, to the best of our knowledge.
@@
Contributions of this article –We characterise these gauge freedoms by establishing necessary and sufficient conditions under which two representations of the quantum master operator lead to the same PDP. This is achieved by analysing the generators of the PDPs, which constitute unravelled QMEs. In contrast to the QME, they govern the dynamics of probability distributions for pure density matrices, which represent conditional system states. In fact, they fully determine the corresponding PDPs, and thus the ensembles of quantum trajectories. Interestingly, the resulting gauge freedoms are much richer than simple multiplication of jump operators by phase factors or even their permutations; this is particularly apparent in systems with reset jumps, which correspond to jump operators of rank 1, meaning that each jump operator resets the system’s conditional state to a fixed destination.
@@
Implications – In the physical context, gauge invariance or equivalence of PDPs and quantum trajectory ensembles are relevant for weak symmetries of quantum master equations, and whether these are inherited by quantum trajectory ensembles. These issues are explored in [52], which relies extensively on the results presented here. When the goal is to generate quantum trajectory ensembles either in experiments or numerical simulations, our work clarifies the remaining gauge freedom that can be used to optimise these setups, e.g., by implementing a minimal equivalent representation. Our results also reinforce the special status of reset jumps, which are common in many physical scenarios; for example energy level transitions in quantum optics experiments [53, 54, 55], superconducting circuits [56, 35] and quantum dots [57, 58]. They also appear in the context of many-body quantum systems, e.g., when global projective measurements are performed at fixed rate [59]. Finally, there has been much recent interest in a notion of ‘resetting’ of quantum systems [60, 61, 62], which when implemented at a constant rate can be encoded into the standard master equation formalism using necessarily global reset jumps.
The paper is structured as follows. In Sec. 2 we review the QME and unravelled QMEs which prescribe stochastic quantum trajectories corresponding to different representations of the former. In Sec. 3 we give the conditions for the unravelled QMEs for different representations to be the same and the resulting implications. Sec. 4 discusses labelled and partially-labelled quantum trajectories, and the equivalence conditions for associated stochastic processes. Sec. 5 presents an illustrative example. Sections 6 and 7 detail the derivation of the results given in the sections 3 and 4 respectively. We conclude in Sec. 8.
2 System State Dynamics
@@
2.1 Quantum master equation
@@
is called a representation of the quantum master operator (which includes $d$ jump operators in this case). It is assumed that all $J_{j}\neq 0$.
@@ 2.2 Quantum trajectories
is the transition rate facilitated by jump operator of type $k$, i.e., the rate for jump of type $k$, from current state $\psi$ into state $\psi^{\prime}$. In Eq. (13), we introduced a gradient, denoted $\nabla$. The action of this gradient on scalar functions $f(\psi)$ gives the matrix with elements $(\nabla f)_{ab}=\partial f/\partial\psi_{ab}$. The divergence of a matrix $M(\psi)$ is therefore ${\nabla\cdot M=\sum_{ab}(\partial M_{ab}/\partial\psi_{ab})}$ [67]. The integration runs over all Hermitian matrices, see Appendix 1 of [66] for details.
@@
where $\rho_{t}$ is the average state of the system (in general a mixed state). Taking the time derivative of this equation and using Eq. (15) with $f(\psi)=\psi$ shows that $\rho_{t}$ indeed obeys the QME in Eq. (1). [^2] This $f(\psi)$ is a matrix valued function; in such cases $\mathcal{W}$ acts separately on each matrix element.
@@
3 Gauge Invariance for Quantum Trajectories
This Section formulates the gauge freedoms of the unravelled quantum dynamics. These gauge freedoms are found by identifying which representations of the QME lead to the same unravelled generator. The results are stated here, with proofs given Sec. 6.
@@ 3.1 SJED
$k,k^{\prime}\in S_{\alpha}\quad\Leftrightarrow\quad\forall{|\psi\rangle}\;\;\exists\,{c,c^{\prime}}\;\;c\,J_{k}|\psi\rangle=c^{\prime}J_{k^{\prime}}|\psi\rangle,\;\;$ (17)
@@
From a physical perspective, reset SJEDs are interesting because rank-1 jump operators appear in many physical settings, as discussed in Sec. 1. Furthermore, for dynamics with jump operators of this type only, a semi-Markov mapping of the unravelled dynamics exists [66, 67, 68], which simplifies the sampled space in stochastic simulations. Non-reset SJEDs can be relevant in experimental settings if the environment consists of multiple identical reservoirs or classical noise is present in the measurement process. In simulations it would be natural to exploit the associated gauge freedom and combine them into a single jump operator from the start.
3.2 Equality of unravelled generators
@@
the corresponding SJEDs are denoted as $S_{1},\dots,S_{d_{C}}$ and $\tilde{S}_{1},\dots,\tilde{S}_{\tilde{d}_{C}}$, and give rise to the associated super-operators for the composite action of their jump operators [cf. Eq. (18)]
@@
The numbers of SJEDs are $d_{C}\leq d$ and $\tilde{d}_{C}\leq\tilde{d}$. Then the conditions for these two representations to describe the same PDP are given by the following Theorem.
@@
$\tilde{d}_{C}=d_{C}\quad\text{and}\quad\tilde{\mathcal{A}}_{\alpha}=\mathcal{A}_{\pi_{\textsc{c}}(\alpha)}\;\;\forall\alpha,\;$ (28b)
for some permutation $\pi_{\textsc{c}}$ of $\{1,2,\dots,d_{C}\}$.
Note that the definition of SJED ensures that the permutation $\pi$ appearing in Theorem $1$ is uniquely defined (it is not possible that ${\cal A}_{\alpha}={\cal A}_{\beta}$ for $\alpha\neq\beta$). The fact that Eq. (28) is sufficient for Eq. (27) can be verified directly from the definition of ${\cal W}$, with the aid of Eq. (20), as we now explain. Showing the converse requires more work, this proof is given in Sec. 6. The proof shows that the algebraic condition on jump operators Eq. (28b) is sufficient to ensure that the stochastic jumps of the two representations occur between the same quantum states, with the same rates. This result has potential relevance more generally in stochastic processes, beyond the specific case of quantum trajectories.
@@
This case is considered explicitly in the proof in Sec. 6. In general, while Eq. (28b) constrains the composite action of SJEDs, it leaves more freedom in the choice of jump operators, compared with Eq. (29) (see Sec. 3.3, below).
An important corollary of Theorem 1 is that $\mathcal{W}^{\dagger}=\tilde{\mathcal{W}}^{\dagger}$ under the conditions in Eq. (28), which means that their stationary distributions over conditional states are identical, and also $P(\psi,t)$ at any time $t$ is the same for both representations provided that one considers the same initial distribution. The two different representations actually produce the same ensemble of quantum trajectories in that case [but Eq. (28) does not guarantee equivalence of their measurement records, see Sec. 4]. We clarify what transformations between representations allow for that invariance next.
@@
3.3 Gauge freedoms of unravelled quantum dynamics
Given Theorem 1, a natural question follows: which representations give rise to a given a set of super-operators $\{{\cal A_{\alpha}}\}_{\alpha=1}^{d_{C}}$, and hence to the same unravelled dynamics? To answer this, we note that equality of $\tilde{\mathcal{A}}_{\alpha}$ and $\mathcal{A}_{\pi_{\textsc{c}}(\alpha)}$ in Eq. (28b) is the condition [cf. Eq. (18)]
$\sum_{j\in\tilde{S}_{\alpha}}\tilde{\mathcal{J}}_{j}=\sum_{k\in S_{\pi_{\textsc{c}}(\alpha)}}{\mathcal{J}}_{k}\,.$ (30)
From Eq. (8), it then follows that the sets $\{\tilde{J}_{j}\}_{j\in\tilde{S}_{\alpha}}$ and $\{J_{k}\}_{k\in S_{\pi_{\textsc{c}}(\alpha)}}$ are different representations of the same completely positive super-operator, $\tilde{\mathcal{A}}_{\alpha}=\mathcal{A}_{\pi_{\textsc{c}}(\alpha)}$, and thus are related by an isometry [64]. Appendix A.2 describes the gauge freedoms associated with the composite action operator for each SJED.
Using these results, we now describe the full gauge freedom of the unravelled quantum dynamics, see Appendix A.3 for details. It is convenient to consider a minimal representation (in the sense that each ${\cal A}_{\alpha}$ is represented by a minimal number of jump operators, see Appendix A). Then the gauge freedoms of the unravelled dynamics are given as follows:
Suppose that $H^{\prime},J_{1}^{\prime},\dots,J_{d^{\prime}}^{\prime}$ is a representation of a given quantum master operator in which all SJED actions ${\cal A}^{\prime}_{1},\dots,{\cal A}^{\prime}_{d_{C}^{\prime}}$ have minimal representations. Then $H,J_{1},\dots,J_{d}$ has the same unravelled generator if and only if $d_{C}^{\prime}=d_{C}$ and there exists a permutation $\pi_{\textsc{c}}$ of $\{1,\dots,d_{C}\}$ and a matrix $\mathbf{V}\in\mathbb{C}^{d\times d^{\prime}}$ such that
@@ 3.3 Gauge freedoms of unravelled quantum dynamics
$\mathbf{V}_{jk}^{(\alpha)}=0\quad\text{unless}\quad j\in S_{\alpha},\,k\in S^{\prime}_{\pi_{\textsc{c}}(\alpha)}$ (31c)
@@
$\sum_{j\in S_{\alpha}}[\mathbf{V}^{(\alpha)}_{jk}]^{*}\mathbf{V}^{(\alpha)}_{jk^{\prime}}=\delta_{kk^{\prime}}\quad\text{for}\quad k,k^{\prime}\in S^{\prime}_{\pi_{\textsc{c}}(\alpha)}\;.$ (31d)